A car traveles at a constant speed of 37km/hr round of curve of radius 300m find how long it will take to turn through 180 degree of Arc.


1
Expert's answer
2021-11-14T17:14:14-0500
"T=\frac{\pi R}{v}\\T=\frac{\pi (300)}{37}3.6=91.7\ s"